Company X uses a job cost system and applies overhead to production on the basis of direct labor cost. On January 1, 2018, Job 50 was the only job in process. The costs incurred prior to January 1 on this job were as follows: DM of TL27,000, DL of TL11,600, and MOH of TL15,008. As of January 1, Job 49 had been completed at a cost of TL85,004 and was part of finished goods inventory. There was a TL13,500 balance in the raw materials inventory account. During the month of January, Company X began production on Jobs 51 and 52, and completed Jobs 50 and 51. Jobs 49 and 50 were sold on account during the month for TL122,600 and TL158,040, respectively. Following additional events occurred during the month:

• Purchased additional raw materials of TL95,007 on account.

• Incurred factory labor costs of TL64,600 exclusive of payroll taxes. Payroll taxes amounted to TL16,080.

• Incurred indirect materials of TL17,000, and indirect labor of TL20,008

• Depreciation on factory equipment totalled to TL12,000

• Other factory expenses on account were TL14,000 2

• During the month, Company X assigned following costs to respective jobs as follows:

Assigned Costs DM DL

Job 50 10,000 5,700

Job 51 39,000 25,040

Job 52 30,800 20,000

Company X estimates total manufacturing overhead costs of TL887,000, direct labor costs of TL700,080 and direct labor hours 20,060 for the year.

Q-1) Compute the balance of WIP as at January 31.

Q-2) Compute the balance of FG as at January 31.

Q-3) Calculate the balance of MOH account as at January 31.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Company X
Predetermined overhead rate Amount TL Note
Expected MOH costs 887,000.00 A
Estimated direct labor costs 700,080.00 B
Predetermined overhead rate 126.70% C=A/B
Job Cost sheet
Particulars Job 50 Job 51 Job 52 Total Note
Opening work in process
Direct Materials      22,700.00     22,700.00
Direct Labor      11,600.00     11,600.00
Overhead      15,008.00     15,008.00
Cost incurred during the month
Direct Materials      10,000.00 39,000.00     30,800.00     79,800.00
Auditor Labor costs        5,700.00 25,040.00     20,000.00     50,740.00 D
Overhead applied        7,221.89 31,725.63     25,339.96     64,287.48 E=C*D
Total Cost     72,229.89 95,765.63     76,139.96 244,135.48 F
Answer 1
Job 52
Work in Process Inventory $ 76,139.96
Answer 2 Job 51
Finished goods Inventory $ 95,765.63
Answer 3
Balance of MOH account
Amount TL
Indirect materials      17,000.00
Indirect labor      20,008.00
Depreciation on factory equipment      12,000.00
Other factory expenses      14,000.00
Actual overhead     63,008.00 G
Less: Overhead applied
     64,287.48
See E
Balance of MOH account (Credit balance)        1,279.48 H=E-G